(3) [1 mark] Find the step response of the unity negative feedback system with the forward path trasfer function of

Without the use of matlab.
Answer #1

R(s) sti cts) gain of the system stlal sti S +2 Given ret)=1 Pils) = Y cl) = 5 ste (69) = 51642)c(5)= S+2 31642) - test S+ 2 I- Als+2) + BS Compare s term Coefficient 0 = A+Bo A = - 80 Compare constant term Coefficient T=